OpenStudy (studygirl303):

Help pleasee!! b/12=16

OpenStudy (misssmartiez):

Are we finding b?

OpenStudy (studygirl303):

Yes.

OpenStudy (misssmartiez):

12 x 16 = b. Basically, to find it isolated variable, you do inverse operations.

OpenStudy (misssmartiez):

12 x 16 = 192 192/12 = 16

OpenStudy (studygirl303):

So im just multiplying 12 by 16 to get my answer?

OpenStudy (misssmartiez):

Yes!

OpenStudy (studygirl303):

Thanks!!

OpenStudy (misssmartiez):

Examples: 5 + v = 12. 12 - 5 = v, which is 7. We take the known variables and cope with those.